# cryoSPHERE: Single-particle heterogeneous reconstruction from cryo EM

### v1

@@ A.2 Molecular dynamics dataset
Figure 13: Left: comparison of a ground truth volume and the corresponding reconstructed volume from cryoSPHERE. Gray is cryoSPHERE, purple is the ground truth. Middle left: comparison of a ground truth volume and the corresponding reconstructed volume from cryoDRGN. Pink is cryoDRGN, purple is the ground truth. Middle right: cryoSPHERE. Right: cryoDRGN.
Figure 14: left: comparison of a ground truth volume and the corresponding reconstructed volume from cryoSPHERE with the unfitted AlphaFold structure. Gray is cryoSPHERE, purple is the ground truth. Middle left: comparison of a ground truth volume and the corresponding reconstructed volume from cryoSPHERE with the fitted Alphafols structure. Blue is fitted, purple is the ground truth. Middle right: cryoSPHERE without fitting. Right: cryoSPHERE with fitting.
Figure 15: Comparison of volumes for an image of the MD 150k dataset. Gray is cryoSPHERE with unfitted AlphaFold structure and 6 segments. Tan is the same with 16 segments. Blue is ground truth. Having 16 segments leads to a volume matching the ground truth much better than 6 segments. This is true for medium scale elements throughout the volume, even at the bottom of the protein, which tends to stay still. Have a look at the coil on the middle right of the figure. Having 16 segments allows the method to break this coil into a segment and to adjust its position. Which is not the case with only 6 segments.
Figure 16: Plots for the MD 150k dataset starting from the first MD structure. Left: histograms of the true and predicted distances. Middle:predicted against true distances. Right: mean FSC curves for the AlphaFold base structure and the first MD structure as a base structure.
Figure 17: Plots for the MD 150k dataset with the last MD structure. Left: histograms of the true and predicted distances. Middle: Predicted against true distances. Right: Mean FSC curves for the AlphaFold base structure and the last MD structure as a base structure.
Figure 18: Examples of recovered segments. From left to right: 10k dataset, 6 segments. 150k dataset unfitted AlphaFold structure, 6 segments. 150k dataset with the first MD structure as the base structure, 6 segments. 150k dataset with the last MD structure as the base structure, 6 segments. 150k dataset unfitted AlphaFold structure, 16 segments.
A.2.2 10k dataset
To assess the performance under data conditions, we run our method with only $10^{4}$ images, $N_{\textrm{segm}}=6$ and the same unfitted AlphaFold base structure and settings as in subsection A.2.1. We train our network and cryoDRGN for only 12 hours only.
Figure 19 demonstrates that cryoSPHERE maintains an excellent agreement with the ground truth and is not significantly impacted by the scarcity of data. On the contrary, the FSC curves indicate that the performance of cryoDRGN is affected. This is further confirmed by Figure 20, where cryoSPHERE has roughly the same FSC curves between the 10k and 150k datasets, while cryoDRGN experiences a significant drop in performance. This is likely because the AlphaFold structure provides strong prior knowledge.
Additionaly, we plot the recovered segments in Figure 18. Despite having fewer images, our method succesfully divides the protein into meaningful parts for the given task.
Figure 19: For the MD 10k dataset. Left: Histograms of the true and predicted distances. Middle: Predicted against true distances. Right: mean FSC curves for the cryoSPHERE with the same base AlphaFold structure as the 150k dataset and cryoDRGN ran on this 10k dataset. The blue lines are for cryoSPHERE, the red lines for cryoDRGN.
Figure 20: For the MD 10k dataset. Left: comparison of the FSC curves for cryoSPHERE on the 150k and 10k datasets. Right: comparison of the FSC curves for cryoDRGN on the 150k and 10k datasets. Blue represents the 150k results and red the 10k results.

### v2

@@ B.2 Molecular dynamics dataset
Figure 18: SNR $0.1$. Examples of reconstructed volumes by cryoDRGN. Green is cryoDRGN and gray is the corresponding ground truth volume.
Figure 19: Results for cryoSPHERE on SNR $0.1$ with $N_{\textrm{segm}}=10$. Left: distribution of distances predicted by cryoSPHERE compared to the ground truth distribution. Middle: true versus predicted distances for cryoSPHERE. Right: FSC comparison between cryoSPHERE with $N_{\textrm{segm}}=20$ and $N_{\textrm{segm}}=10$.
B.2.2 SNR $0.01$
This subsection describes the results for cryoSPHERE, cryoStar and cryoDRGN on our molecular dynamics simulation dataset with SNR $0.01$.
Figures 20 and 21 show that both cryoSPHERE and cryoStar are able to recover the ground truth distribution of distance. Also, given an image, both methods are able to recover the correct conformation.
Figure 22 shows that both cryoSPHERE and cryoStar are outperforming cryoDRGN at the $0.5$ and $0.143$ cutoffs while having very similar performances. In addition, we can see that the volume method of cryoStar perform similarly, if not worse, than cryoDRGN. That seems to indicate that this volume method does not benefit from the information gained by the structural method. Figure 23 shows three examples of volumes reconstructed with the volume method of cryoStar, together with the corresponding ground truth.
Figure 24 and 25 shows a set of predicted structures compared to the ground truth. Both cryoSPHERE and cryoStar are able to recover the ground truth almost perfectly.
Figure 20: Results for cryoSPHERE on SNR $0.01$ with $N_{\textrm{segm}}=25$. Left: distribution of distances predicted by cryoSPHERE compared to the ground truth distribution. Middle: true versus predicted distances for cryoSPHERE. Right: segments decomposition.
Figure 21: Results for cryoStar on SNR $0.01$. Left: distribution of distances predicted by cryoStar compared to the ground truth distribution. Right: true versus predicted distances for cryoStar
Figure 22: SNR $0.01$. Mean Fourier shell correlation ($\pm$ std) comparison for cryoSPHERE with $N_{\textrm{segm}}=25$, cryoStar and cryoDRGN. Top left: cryoSPHERE versus cryoDRGN. Top right: cryoSPHERE vs cryoStar. Bottom left: cryoDRGN vs cryoStar. Bottom right: cryoStar volume method vs cryoDRGN.
Figure 23: SNR $0.01$. Example volumes produced by the volume method of cryoStar with the corresponding ground truth. Top: ground truth. Bottom: corresponding volumes predicted by the volume method of cryoStar.
Figure 24: SNR $0.01$. Examples of reconstructed structures by cryoSHERE. Blue is predicted, green is ground truth. cryoSPHERE is able to recover the right conformation. Left to right: image number 11, 5001 and $9$.
Figure 25: SNR $0.01$. Examples of reconstructed structures by cryoStar. Red is predicted, green is ground truth. cryoStar is able to recover the right conformation. Left to right: image number 11, 5001 and $9$.
Figure 26: SNR $0.01$. Examples of reconstructed volumes by cryoDRGN. Green is cryoDRGN and gray is the corresponding ground truth volume.
Finally, Figure 26 shows examples of cryoDRGN predicted volume together with the corresponding ground truth. The method is able to recover the ground truth volumes almost perfectly, with a somewhat lower resolution compared to Figure 18.
B.2.3 SNR $0.001$
This subsection complements the results of Section 5.1.
Figure 27 shows that cryoStar is also able to recover the rough distribution of conformations as well as the correct conformation given an image. The comparisons of FSC shows that cryoStar perform similarly to cryoDRGN at the $0.5$ cutoff but better at the $0.143$. In addition, Figure 28 shows that cryoStar volume method does not perform significantly better than cryoDRGN, in spite of using the information given by the structural method of cryoStar. Figure 31 shows a set of example volumes reconstructed by the volume method of cryoStar and the corresponding ground truth.
Figure 29 and 30 show a set of predicted structures compared to the ground truth. Both cryoSPHERE and cryoStar are able to recover the ground truth, but in a more approximate fashion than for the higher SNR datasets.
Figure 27: Results for cryoStar on SNR $0.001$. Left: distribution of distances predicted by cryoStar compared to the ground truth distribution. Middle: true vs predicted distances for cryoStar. Right: FSC comparison between cryoStar and cryoDRGN
Figure 28: SNR $0.001$. Mean Fourier shell correlation ($\pm$ std) comparison for cryoStar and cryoDRGN. Left: cryoStar versus cryoDRGN. Right: cryoStar volume method vs cryoDRGN.
Figure 29: SNR $0.001$. Examples of reconstructed structures by cryoSHERE. Blue is predicted, green is ground truth. CryoSPHERE is able to recover the right conformation. Left to right: image number 11, 5001 and $9$.
Figure 30: SNR $0.001$. Examples of reconstructed structures by CryoStar. Red is predicted, green is ground truth. CryoStar is able to recover the right conformation. Left to right: image number 11, 5001 and $9$.
Figure 31: SNR $0.001$. Example volumes produced by the volume method of cryoStar with the corresponding ground truth. Top: ground truth. Bottom: corresponding volumes predicted by the volume method of cryoStar.
Figure 32: SNR $0.001$. Examples of reconstructed volumes by cryoDRGN. Green is cryoDRGN and gray is the corresponding ground truth volume.
Figure 32 shows examples of cryoDRGN predicted volume together with the corresponding ground truth. The method underestimates the opening of the protein and predicts low resolution volumes with a lot of noise. It seems it is overfitting.
B.2.4 Debiasing cryoSPHERE
When deforming an atomic model to recover different conformations, one should be careful not to bias the results. CryoStar (Li et al., 2023) developed a volume method to help assess the bias induced by the atomic model. However, we demonstrate here that this two stages of training are nothing specific to cryoStar and can, in fact, be applied to any structural method just by using DRGN-AI Levy et al. (2024).
We remove the last 40 residues of chain B of the 10 000 structures obtained through molecular dynamics simulations, see Appendix B.2. We then follow the exact same process to generate 150k images with SNR 0.01. Again, see Appendix B.2.
We run cryoSPHERE with $N_{\textrm{segm}}=20$ for 24 hours using the same base structure as the other experiments of Appendix B.2, that we obtained through AlphaFold. In other words, our base structure has 40 more residues than the structures on the images, see Figure 33.
Figure 33: Left: example of a structure used to to generate the images. The missing part is highlighted in a red ellipsoid. Right: the base structure used, with the residues that were removed for the image generation highlighted in a black ellipsoid.
Figure 34: Debiasing the atomic model. Left: Distribution of ground truth and predicted distances. Right: True versus predicted distances.
Figure 35: Debiasing the atomic model. Top, from left to right: example of 3 structures predicted by cryoSPHERE. The red ellipses show the alpha helix that is present in the base structure but is not on the images. Bottom, from left to right: corresponding volumes reconstructed by DRGN-AI.
This discrepancy does not affect the predicted distances by cryoSPHERE, as shown in Figure 34. Figure 35 shows 3 structures predicted by cryoSPHERE. Even though the opening motion is correctly recovered, the algorithm tries to remove the missing alpha helix from where there should not be any density. In a sense, cryoSPHERE detects that the atomic model has too many residues, but it cannot temove them, by design. We leave to future work the ability to learn the amplitude of each Gaussian mode of Equation 1. Here, we propose to debias the atomic model by running, in a second step, DRGN-AI with the final latent space of the cryoSPHERE run and fixed poses, similar to what cryoStar proposes. Figure 35 shows the reconstructed volumes by DRGN-AI corresponding to the plotted structures predicted by cryoSPHERE. The alpha helix is not present, which permits to detect the bias brought by the base structure for cryoSPHERE.
In terms of computational cost, on a dataset of 150k images of size $190\times 190$, DRGN-AI performs one epoch in 10 minutes while the volume method of cryoStar perform one epoch in 6 minutes.
B.2.5 Comparison accross SNR for cryoSPHERE, cryoDRGN and cryoStar.
Figure 36: Segmentation of cryoSphere with $N_{\textrm{segm}}=10$. Left: SNR $0.1$. Middle: SNR $0.01$. Right: SNR $0.001$.
Figure 37: Comparison, for each method, of the Fourier shell correlation for different SNR. Left to right: cryoSPHERE, cryoStar, cryoDRGN.
In this section, we look at the change of performance for each method accross the different SNR.
Figure 37 shows that all three methods experience a drop in FSC with decreasing SNR. CryoSPHERE and cryoStar show a drop of one standard deviation between SNR $0.1$ and $0.01$ while cryoStar experience a much bigger drop than cryoSPHERE between SNR $0.01$ and $0.001$, which confirms that cryoSPHERE is more resilient to a high level of noise.
CryoDRGN shows a steady decrease in its FSC with decreasing SNR.
The fact that for a SNR of $0.1$ cryoStar is slightly outperforming cryoSPHERE (see Appendix B.2.1), that for a SNR of $0.01$ cryoSPHERE outperforms cryoStar (see Appendix B.2.2), and that for a SNR of $0.001$ cryoSPHERE outperforms cryoStar by one standard deviation (see Section 5.1) confirms that moving big chunks of the protein as rigid bodies is more resilient to low SNR than moving each residue individually.
B.2.6 Comparison for different values of $N_{\textrm{segm}}$.
In this subsection, we compare the results of cryoSPHERE for three different values of $N_{\textrm{segm}}=10,20,25$, for the different SNR of the MD dataset in 5.1.
Figure 38: FSC curves for different $N_{\textrm{segm}}$ values. From left to right: SNR 0.1, SNR 0.01, SNR 0.001.
Figure 38 shows the evolution of the FSC curves with different values of $N_{\textrm{segm}}$ for different SNR. As we can expect, the high the number of segments, the more flexible cryoSPHERE and the better FSC. The lower the SNR, the greater we gain in FSC by increasing $N_{\textrm{segm}}$. This is because with decreasing SNR, the initial fitting of the structure in a consensus reconstruction is less accurate. Hence, the method benefits from a greater flexibility to adjust the protein on a smaller scale.
Figure 39: MD dataset, SNR 0.1. Distribution of the predicted distances. From left to right: 10 segments, 20 segments, 25 segments.
Figure 40: MD dataset, SNR 0.01. Distribution of the predicted distances. From left to right: 10 segments, 20 segments, 25 segments.
Figure 41: MD dataset, SNR 0.001. Distribution of the predicted distances. From left to right: 10 segments, 20 segments, 25 segments.
Figure 42: MD dataset, SNR 0.1. True versus predicted distances. From left to right: 10 segments, 20 segments, 25 segments.
Figure 43: MD dataset, SNR 0.01. True versus predicted distances. From left to right: 10 segments, 20 segments, 25 segments.
Figure 44: MD dataset, SNR 0.001. True versus predicted distances. From left to right: 10 segments, 20 segments, 25 segments.
Figures 39, 40 41,42,43,44 show the distributions of the predicted and true distances and the true versus the predicted distances for each value of SNR and $N_{\textrm{segm}}$. This shows that the choice of $N_{\textrm{segm}}$ is not so critical for cryoSPHERE to work well. The higher $N_{\textrm{segm}}$, the better it is in terms of FSC, but a value of 10 still gives a good performance.
Figure 45: MD dataset, SNR 0.001. Segments decomposition. From left to right: 10 segments, 20 segments, 25 segments.
Figure 46: MD dataset, SNR 0.01. Segments decomposition. From left to right: 10 segments, 20 segments, 25 segments.
Figure 47: MD dataset, SNR 0.1. Segments decomposition. From left to right: 10 segments, 20 segments, 25 segments.
Finally, Figures 45,46and 47 show the segment decomposition for different values of $N_{\textrm{segm}}$. CryoSPHERE always learns a segmentation differentiating the two chains and the top and bottom parts of the protein.
B.3 EMPIAR-10180
This section gives more details on the experiment with the EMPIAR-10180 datasets described in Section 5.2. The data was processed by Relion, hence the poses and CTF are assumed to be known. We use an encoder with 4 hidden layers of size 512, 256, 64, 64 and a decoder with 2 hidden layers with size 512, 512.
We train cryoSPHERE with no clashing nor continuity loss, with $N_{\textrm{segm}}=20$. We use the ADAM optimizer with a learning rate of 0.00003 for the parameters of the encoder and decoder, while we set the learning rate to 0.0003 for the parameters of the GMM segmentation.
We low pass filter the images with a bandwith of $23.4$Å, we apply a mask of radius $0.9375$ to the input images and we apply a mask of radius $1$ to the true and predicted images for the computation of the correlation loss.
Figure 49 shows 4 structures taken from the principal component traversal depicted in Figure 48. The structures contain only the $C_{\alpha}$ atoms. We provide a movie of the traversals of principal component 1 and principal component 2 by clicking here.
We subsequently train DRGN-AI on the latent space provided by cryoSPHERE, similar to cryoStar Phase II of training. We show volumes taken along the first principal component in Figure 50. We recover the correct bending of the protein toward its ”foot”. In addition, this second step detects the compositional heterogeneity and the density is zero in this region. This is a detail that the structural method of cryoSPHERE could not detect. We successfully identified a bias in the base structure.
We provide a movie of the motion recovered by DRGN-AI trained on the latent variable of cryoSPHERE here.
Figure 48: Empiar 10180: kernel density plot of the first and second principal components of that latent space of cryoSPHERE. The red dots are the point selected for the traversal of the first principal component.
Figure 49: Empiar 10180. Four structures taken along the first principal component, from blue to white to red. Top: view from the ”back”. Bottom: view from the ”top”.
Figure 50: Empiar 10180, DRGN-AI is trained on the latent space of cryoSPHERE. Three Volumes taken evenly along the principal component.
B.4 EMPIAR-12093
This experiment demonstrates that cryoSPHERE is applicable to real data with high noise levels. We applied cryoDRGN, cryoSTAR and cryoSPHERE to a bacterial phytochrome (Bódizs et al., 2024) dataset (medium-sized protein,  120 kDa). This dataset comprises two distinct subsets of 200 000 images of size 400 x 400 each, representing the protein in its light-activated state (red-absorbing state, called Pr) and its resting state (far-red-absorbing state, called Pfr). The pre-processing steps are detailed in Bódizs et al. (2024). We downgrade the images to a size of 256 x 256. We give a computational budget of 24 hours on the same single GPU to cryoDRGN, cryoStar structural method and cryoSphere and compare their results. We subsequently run cryoStar volume method on the latent space obtained by cryoStar and cryoStar volume method on the latent space obtained by cryoSPHERE.
We provide PC 1 traversal movies for both cryoSPHERE and cryoStar structures and volumes for both Pr and Pfr here.
B.4.1 Pfr state
We provide the first PC traversal for cryoDRGN in Figure 51. CryoDRGN does not recover the upper part of the protein at all. In addition, there is no motion through the principal component. This might indicate that the recovered motion is in fact noise in the top part. We additionally plot 3 structures taken evenly along the first principal component of the cryoStar volume method in Figure 52. The method is also unable to reconstruct the very mobile top part, in spite of using the latent space of the structural method. For Pfr, the debiasing technique of cryoStar through a volume method is ineffective for the top part.
Finally, we run cryoStar volume method on the latent space recovered by cryoSPHERE and show the first principal component traversal in Figure 53. Similar to cryoDRGN and cryoStar Phase II, this procedure is unable to reconstruct the top part of the protein. Hence this allows debiasing on the bottom part of the protein only.
We provide a movie of the first PC traversal of cryoStar volume method with cryoSPHERE latent variables here and the same movie with cryoStar latent variables here.
Figure 51: Pfr: 3 volumes taken evenly along the first principal component of cryoDRGN.
Figure 52: Pfr: 3 volumes taken evenly along the first principal component of cryoStar volume method.
Figure 53: Pfr state. CryoStar volume method is trained on the latent space of cryoSPHERE. From left to right: three volumes taken evenly along the first principal component.
B.4.2 Pr state
We provide three structures taken evenly along the first principal component of cryoDRGN in Figure 54. There is no motion in the bottom part as expected. However, the method is unable to recover the top part.
We also run cryoStar volume method on the latent space recovered by cryoSPHERE and display the first principal component traversal in Figure 56. We encounter the same difficulties as cryoStar volume method and cryoDRGN, see Figures 54, 55. CryoDRGN is not able to recover the top part of the protein. While cryoStar volume method does recover some motion for this top part, the resolution is too low to debias the base structure. Therefore, this debiasing procedure is only useful for the lower part of the protein.
We provide a movie of the first PC traversal of cryoStar volume method with cryoSPHERE latent variables here and the same movie with cryoStar latent variables here.
Figure 54: Pr: 3 volumes taken evenly along the first principal component of cryoDRGN volume method. The top and bottom volumes are the same with a different density threshold.
Figure 55: Pr: 3 volumes taken evenly along the first principal component of cryoStar volume method.
Figure 56: Pr state. CryoStar volume method is trained using the latent space of cryoSPHERE. From left to right: volumes taken evenly along the first principal component.
B.5 Computational costs
CryoStar and cryoSPHERE share the same way of turning a structure into a volume. This is a computationally expensive procedure, reduced by the sperability of Gaussian kernels, see e.g (Chen et al., 2023).For an of size $N_{\textrm{pix}}\times N_{\textrm{pix}}$, it involves computing two times the distance of each residues to $N_{\textrm{pix}}$ pixels and taking the product of these vectors to obtain a matrix representing the images for each residue, then summing over the residues. This is one of the computational bottlenecks of the structural methods. In spite of sharing the same bottleneck, cryoSPHERE tends to be slightly more computationally demanding. This is because cryoSPHERE needs to compose $N_{\textrm{segm}}$ rotations for each residue, while cryoStar only translates each residue. For example, for the experiments of Section 5.3, in 24 hours, cryoStar performs roughly 190 epochs and cryoSPHERE with $N_{\textrm{segm}}=20$ performs roughly 130 epochs. Similarly, for the experiment of Section 5.2, cryoStar performs 184 epochs while cryoSPHERE with $N_{\textrm{segm}}$ performs 95 epochs. However, owing to the reduced number of freedom of cryoSPHERE compared to cryoStar, we observe that cryoSPHERE performs as well or better as cryoStar for the same computational budget.